OpenStudy (anonymous):

When you simplify each side of the equation 5n – 2(3 – n) = 4(2n – 3), you will get the equivalent equation 7n – 6 = 8n - 12

OpenStudy (anonymous):

yes ! 5n-2(3-n)=4(2n-3) distribute (-2) to (3) and (-n) 5n-6+2n=4(2n-3) do the same to the other side distribute (4) to (2n) and (-3) 5n-6+2n=8n-12 simplify 7n-6=8n-12

OpenStudy (anonymous):

you forgot 5n^2

OpenStudy (anonymous):

because when (n) (n) = n^2

OpenStudy (anonymous):

wait was the equation supposed to be 5n^2 – 2(3 – n) = 4(2n – 3)???

OpenStudy (anonymous):

no don't change the problem itself, just make sure when you times 5n times n = 5n^2

OpenStudy (anonymous):

-5n^2+15-6+2n=8n-12n

OpenStudy (anonymous):

no thats wrong pemdas states that you would do the multiplication -2(3-n) and then subtract that answer from 5n which would be 5n-6+2n...... you would NOT (5n-2)(3-n) that would be a whole other equation

OpenStudy (anonymous):

so therefor 5n^2+23=8n-12n-2n =-4n-2n=-6n

OpenStudy (anonymous):

no, when you take 5n-2 times 3-n the five gets times like this 5n(3) and (5n)(-n)

OpenStudy (anonymous):

whats is (n) (n)=